John crewing agency has an average of 200 applicants per week for the position of deck cadet with a standard deviation of 12. A random sample of size 4 weeks will be taken.
A) determine the mean and standard error of the randomly selected weeks.
B) find the percentage of the randomly selected weeks to have a sample mean more than 210 applicants.
C) find the percentage of the randomly selected weeks to have a sample mean less than 185 applicants.
D) find the percentage of the randomly selected weeks to have a sample mean within 8 applicants of the population mean.
1
Expert's answer
2020-12-14T10:29:15-0500

"\mu=200"

"\sigma=12"

"n =4"


"\mu_4=\mu\cdot n=800"

"\sigma_4=\sigma\cdot n=48"

"SE=\frac{\sigma_4}{\sqrt n}=\frac{48}{\sqrt4}=24"


More than 210:

"P(X>210)=P(Z>\frac{210-200}{12})=P(Z>0.83)=1-P(Z<0.83)="

"=1-0.7967=0.2033"

Less than 185:

"P(X<185)=P(Z<\frac{185-200}{12})=P(Z<-1.25)=0.1056"

Within 8:

"P(192<X<208)=P(\frac{192-200}{12}<Z<\frac{208-200}{12})=P(-0.67<Z<0.67)="

"=P(Z<0.67)-P(Z<-0.67)=0.7486-0.2514=0.4972"